An ice skater skating along an icy surface starts to coast (stops applying a force to the ice). The 

skater coasts a distance of 95 m in 15.8 s. If the coefficient of kinetic friction between the ice 

and skates was 0.077, what was the initial velocity of the skater? 


1
Expert's answer
2021-05-13T16:23:00-0400

"a=F/m=\mu mg/m=\mu g"


"s=v_0t-at^2/2=v_0t-\mu gt^2/2\to"


"v_0=(s+\mu gt^2/2)/t=(95+0.077\cdot9.81\cdot 15.8^2/2)/15.8=11.98\ (m/s)" . Answer